OpenStudy (anonymous):

parametric equations - find the arc length of x=t^2 , y=4t^(3)-1 on the interval [-1,1] please help

OpenStudy (anonymous):

Sqrt[(dy/dt)^2 + (dx/dt)^2] integrate from -1 to 1

OpenStudy (anonymous):

I keep getting zero and that isn't correct...

OpenStudy (anonymous):

\[\int_{-1}^1 \sqrt{4 t^2+\left(12 t^2\right)^2} \, dt\]

OpenStudy (anonymous):

\[\frac{1}{27} \left(-1+37 \sqrt{37}\right)\]

OpenStudy (anonymous):

I dont understand how you are getting that.....

OpenStudy (anonymous):

what part did you get stuck?

OpenStudy (anonymous):

how are you doing the u substitution?